OpenStudy (miss.rose):

How many quarter notes would have the same duration as 4 eighth notes? Represent this as a fraction. Help Please! I will give medal! @zoeeh127

OpenStudy (ghostgate):

Alright a quarter is 1/4ths of something, just like a quarter is $0.25 cents which is 1/4ths of a dollar. An eighth note is half of a quarter basically, equivalent to 0.125, so... 4 eighth notes would equal 4/8 which is 1/2, well 2 quarter notes equal 2/4 which is 1/2. So your answer is that 2 quarter notes equal 4 eighth notes, so: 2/4 = 4/8 ~ Both equal 1/2. I hope this helps in someway! Have a great day! {---Ghostgate---}

OpenStudy (miss.rose):

Thank you! This has helped me a lot. I will give you your medal. Do you mind helping me with some other problems?

OpenStudy (ghostgate):

Sure, what's the problem?

OpenStudy (miss.rose):

Rewrite each of the following durations using sixteenth notes. Write in words and as a fraction: 3 quarter notes 5 eighth notes 1 half note

OpenStudy (ghostgate):

Okay, based off of what this sounds like, you have a denominator of '16ths'. So... as I mentioned before if a quarter is 1/4ths of something, what is 1/4ths of 16?

OpenStudy (zoeeh127):

i will help

OpenStudy (miss.rose):

4?

OpenStudy (ghostgate):

Yes, so 3 quarters = ?/16ths? Same goes for the others, as I said 1/8ths is half of 1/4ths, so 2 is half of 4 so...: 5 eighth notes = ?/16ths? And half of it should be easy: Half = ?/16ths.

OpenStudy (zoeeh127):

3 quaurter notes is12 sixteenth notes

OpenStudy (miss.rose):

Wait... I am getting confused is it 4/16 or 12/16?

OpenStudy (ghostgate):

3 quarters is 12/16ths since you have 3 4/16ths. Eighth notes are 2/16ths, so 5 of those is 10/16ths. And a half is 8/16ths. Do you get that?

OpenStudy (miss.rose):

Yes, I get that. Can I ask you two more?

OpenStudy (ghostgate):

Sure, go ahead.

OpenStudy (miss.rose):

Okay. Here they are: Rewrite each of the following durations using eighth notes. Write in words and as a fraction: 1 quarter note 1 half note 1 full note Draw using the note symbols 2 different ways to show a duration of 3 eighth notes.

OpenStudy (ghostgate):

1 quarter is 1/4, using the '8th' system it is: 1 quarter = 2/8 1 half = 1/2, using the '8th' system it is: 1 half = 4/8 1 full = 1/1, using the '8th' system it is: 1 full = 8/8 I don't know exactly how to draw them, so... I don't think I can help you on that last part.

OpenStudy (miss.rose):

No problem! Thank you soooooo much for your help! Also, I attached a picture of all the music notes. So take a look at that and tell me if that helps.

OpenStudy (ghostgate):

Okay, I guess to get the equivalent of 3 1/8th notes, you would do this: